\begin{thebibliography}{10}

\bibitem{alvarado2021robust}
Alejandra Alvarado, Angelos Koutsianas, Beth Malmskog, Christopher Rasmussen, Christelle Vincent, and Mckenzie West.
\newblock A robust implementation for solving the {S}-unit equation and several applications.
\newblock In {\em Arithmetic geometry, number theory, and computation}, pages 1--41. Springer, 2021.

\bibitem{anni2016modular}
Samuele Anni and Samir Siksek.
\newblock Modular elliptic curves over real abelian fields and the generalized {Fermat} equation {{\(x^{2\ell}+y^{2m}=z^p\)}}.
\newblock {\em Algebra Number Theory}, 10(6):1147--1172, 2016.

\bibitem{chen2024fermat}
Imin Chen, Aisosa Efemwonkieke, and David Sun.
\newblock Fermat’s last theorem over $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{17})$.
\newblock {\em Canadian Journal of Mathematics}, 76(1):18--38, 2024.

\bibitem{freitas2020asymptotic}
Nuno Freitas, Alain Kraus, and Samir Siksek.
\newblock On asymptotic {Fermat} over $\mathbb{Z}_p$-extensions of $\mathbb{Q}$.
\newblock {\em Algebra \& Number Theory}, 14(9):2571--2574, 2020.

\bibitem{freitas2015elliptic}
Nuno Freitas, Bao~V. Le~Hung, and Samir Siksek.
\newblock Elliptic curves over real quadratic fields are modular.
\newblock {\em Invent. Math.}, 201(1):159--206, 2015.

\bibitem{freitas2015asymptotic}
Nuno Freitas and Samir Siksek.
\newblock The asymptotic {Fermat}'s last theorem for five-sixths of real quadratic fields.
\newblock {\em Compos. Math.}, 151(8):1395--1415, 2015.

\bibitem{freitas2015criteria}
Nuno Freitas and Samir Siksek.
\newblock Criteria for irreducibility of mod {{\(p\)}} representations of {Frey} curves.
\newblock {\em J. Th{\'e}or. Nombres Bordx.}, 27(1):67--76, 2015.

\bibitem{gpvillagra}
Franco Golfieri, Ariel Pacetti, and Franco Villagra, Lucas.
\newblock \textquotedblleft on endomorphism algebras of $\text{GL}_2$-type abelian varieties and diophantine applications''.
\newblock To be appeared in Revista Matemática Iberoamericana.

\bibitem{HORIE_2002}
Kuniaki Horie.
\newblock Ideal class groups of iwasawa-theoretical abelian extensions over the rational field.
\newblock {\em Journal of the London Mathematical Society}, 66(2):257–275, 2002.

\bibitem{lang2012cyclotomic}
Serge Lang.
\newblock {\em Cyclotomic fields I and II}, volume 121.
\newblock Springer Science \& Business Media, 2012.

\bibitem{massoutie1931dernier}
L.~Massoutié.
\newblock Sur le dernier th{\'e}or{\`e}me de {Fermat}.
\newblock {\em C. R. Acad. Sci., Paris}, 193:502--504, 1931.

\bibitem{mccall1995imaginary}
Thomas~M. McCall, Charles~J. Parry, and Ramona~R. Ranalli.
\newblock Imaginary bicyclic biquadratic fields with cyclic 2-class group.
\newblock {\em J. Number Theory}, 53(1):88--99, 1995.

\bibitem{pomey1931nouvelles}
Leon Pomey.
\newblock Nouvelles remarques r{\'e}latives au dernier th{\'e}or{\`e}me de {Fermat}.
\newblock {\em C. R. Acad. Sci., Paris}, 193:563--564, 1931.

\bibitem{Serre87}
Jean-Pierre Serre.
\newblock Sur les repr{\'e}sentations modulaires de degr{\'e} 2 de {{\(\text{Gal}({\overline {\mathbb Q}}/{\mathbb Q})\)}}.
\newblock {\em Duke Math. J.}, 54:179--230, 1987.

\bibitem{siksek2024curves}
Samir Siksek and Robin Visser.
\newblock Curves with few bad primes over cyclotomic $\mathbb{Z}_\ell$ extensions.
\newblock {\em Algebra \& Number Theory}, 2024.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage[utf8]{inputenc}
\usepackage[active]{srcltx}
\usepackage{tabularx}
\usepackage{mathrsfs}
\usepackage{array,booktabs,arydshln,xcolor}
\usepackage{fullpage}
\usepackage{amsfonts,enumerate}
\usepackage[mathscr]{euscript}
\usepackage{epsfig}
\usepackage{latexsym}
\usepackage{leftindex}
\usepackage[all]{xy}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{float}
\usepackage{amsmath}
\usepackage{amsxtra}
\usepackage{xcolor}
\usepackage[colorlinks]{hyperref}
\usepackage{fancyvrb}
\usepackage{tikz}
\hypersetup{citecolor=blue}
\usepackage{caption} 
.tfm
.tfm

\theoremstyle{plain}
\newtheorem{prop}{Proposition}[section]
\newtheorem{thm}[prop]{Theorem}
\newtheorem*{thm*}{Theorem}
\newtheorem{coro}[prop]{Corollary}

\newtheorem{lemma}[prop]{Lemma}

\newtheorem{ass}{Assumption}
\newtheorem*{lemma*}{Lemma}
\newtheorem*{prop*}{Proposition}

\theoremstyle{definition}

\theoremstyle{remark}

\newtheorem{remark}{Remark}

\numberwithin{table}{section}

\newcommand{\GL}{{\rm GL}}

\makeatletter
\newcommand{\verbatimfont}[1]{\renewcommand{\verbatim@font}{\ttfamily#1}}
\makeatother

\def\ZZ{\mathbb Z}

\def\FF{\mathbb F}
\def\QQ{\mathbb Q}

\def\ff{\mathfrak f}
\def\pp{\mathfrak p}

\def\qq{\mathfrak q}

\def\<#1>{{\left\langle{#1}\right\rangle}}

\usetikzlibrary{calc}

\begin{document}
	
	\title{On Fermat's Last Theorem over the $\mathbb{Z}_3$-extension of $\mathbb{Q}$ and other fields}
	
	\author{Luis Dieulefait}
	\address{Universitat de Barcelona, G. V. de les Corts Catalanes 585, 08007 Barcelona, Spain} 
	\email{ldieulefait@ub.edu}
	\author{Franco Golfieri Madriaga}
	\address{Center for Research and Development in Mathematics and Applications (CIDMA),
		Department of Mathematics, University of Aveiro, 3810-193 Aveiro, Portugal}
	\email{francogolfieri@ua.pt}

	\keywords{}

		\begin{abstract}

            The main result of the present article is a proof of Fermat's Last Theorem for sufficiently large prime exponents $p$ with $p \equiv 2 \pmod{3}$ over certain number fields. A particular case of these fields are the maximal real subfields of the cyclotomic extensions $\QQ(\zeta_{3^n})$ for every $n$. 
            Our strategy consists in combining the modular method with a generalization of an arithmetic result of Pomey to these fields.
            
		\end{abstract}
	
	\maketitle

	\section{Introduction}

	
   
    Let $K$ be a number field and let denote $\mathcal{O}_K$ its ring of integers. Let us consider the Fermat equation 
    \begin{equation}
    \label{fermat:eq}
    x^p+y^p=z^p, \ \ x,y,z \in \mathcal{O}_K.
    \end{equation}
    We say that the \textbf{asymptotic Fermat's Last Theorem} holds for $K$ if there exists a constant $C_K$, depending only on $K$, such that for all primes $p>C_K$, the only solutions $(a,b,c) \in \mathcal{O}_K^3$ to the equation \eqref{fermat:eq} are the trivial ones, i.e. the solutions satisfying $abc=0$. In \cite{freitas2015asymptotic} and \cite{freitas2015elliptic}, Freitas et al. proved the asymptotic Fermat Last's Theorem over totally real quadratic number fields. In their work, they proved it by studying the $S$-unit solutions to the unit equation. 
    
    Recently, motivated by the work of Freitas, Siksek, and Le Hung, there has been a great advance in the study of asymptotic Fermat's Last Theorem. In this article, we present alternative conditions for our fields $K$, independent from the conditions of the $S$-units stated by Freitas and Siksek (see \cite[Theorem 3]{freitas2015asymptotic}).
  
  In 1931, Massouti\'e \cite{massoutie1931dernier} and Pomey \cite{pomey1931nouvelles} proved the following result:
    
    \begin{thm*}
    	\label{thm:pomey}
    If $p$ is an odd prime such that $p \equiv 2 \pmod 3$ and $x,y,z$ are non-zero integers such that $x^p+y^p+z^p=0$, then $3$ divides $x,y,$ or $z$.
    \end{thm*}

    In their proof, they use the following elementary result.
    \begin{lemma*}
    Let $p$ be a rational prime and $a,b$ rational coprime integers. If $p|a^2+3b^2$, then $p$ is of the form $x^2+3b^2$. 
    \end{lemma*}
    This Lemma is the main obstacle to generalizing Pomey's result for general number fields. Motivated by Pomey's proof, we can generalize the above lemma by adding an extra hypothesis to the class field of $K$ and the ramification of $3$.
    In order to state the result, we need to consider the following notation. Let $K$ be a totally real number field and let $q$ be an odd prime. Let us denote,
    \begin{equation}\label{eq:first_conditions}
    \begin{split}
    S_q &= \{\qq : \qq \ \text{is a prime of } K \ \text{dividing} \ q\},\\
    T_q = \{\qq \in &S_q : f(\qq|q) = 1\}, \ \ V_q= \{\qq \in S_q : 2 \nmid e(\qq|q) \},
    \end{split}
    \end{equation}
    where $f(\qq|q)$ denotes the residual degree of $\qq$. 
    
    We also make the following assumption.  
\begin{ass}
Keeping the previous notation we say that $K$ satisfies Assumption~\ref{assumption} if
\begin{enumerate}
\item $T_3 \neq \emptyset$,
\item $V_3 \neq \emptyset$,
\item $h(K(\sqrt{-3}))$ is odd.
\end{enumerate}
\label{assumption}  
\end{ass}

Then, in section \ref{3nmidabc} we are able to prove the following  theorem.

\begin{thm}
\label{divpomey}
Let $K$ be a totally real field satisfying Assumption~\ref{assumption}. Let $a,b,d \in \mathcal{O}_K$ with $d$ totally positive and let $t:=h(K(\sqrt{-3}))$. Then, if $d|a^2+3b^2$, $d^t$ is of the form $x^2+3y^2$ where $x,y \in \mathcal{O}_K$.
\end{thm}

Motivated by Pomey's proof, Theorem \ref{divpomey} enables us to prove the following generalization of Pomey's theorem for fields satisfying Assumption~\ref{assumption}.

\begin{thm}
\label{pomey_case}
Let $K$ be a totally real field satisfying Assumption~\ref{assumption}. If $p$ is and odd prime such that $p \equiv 2 \pmod 3$ and $(x,y,z) \in \mathcal{O}_K^3$ are non-zero integers such that $x^p+y^p+z^p=0$, then $3$ divides $x,y,$ or $z$.
\end{thm}

Observe that for these particular fields, we cover the case of the asymptotic Fermat's Last Theorem where $3 \nmid abc$, with $(a,b,c)$ a solution to the equation \ref{fermat:eq} (under the assumption that $p \equiv 2 \pmod{3}$). 

To deal with the case $3|abc$ we use modularity results over these number fields and lowering the level results at the prime $3$. To guarantee  modularity, we need the exponent $p$ to be greater than a certain bound that will only depend on $K$.

\begin{thm}
    \label{3midabc}
    Let $K$ be a totally real field. There exists a bound $C_K$, only depending on $K$, such that if $p>C_K$, there are no non-trivial primitive solutions $(x,y,z) \in \mathcal{O}_K^3$ on $K$ to the equation $x^p+y^p+z^p=0$ satisfying $3|xyz$.
    \end{thm}

Finally, Theorem \ref{pomey_case} and Theorem $\ref{3midabc}$ imply our main result.

   \begin{thm}
    \label{thm:main}
  	Let $K$ be a totally real field satisfying Assumption~\ref{assumption}. Then there exists a bound $C_K$ such that if $p \equiv 2 \pmod 3$ and $p>C_K$, there are no non-trivial primitive solutions on $K$ to the equation $x^p+y^p+z^p=0$.
   \end{thm}

   \begin{remark}
   We cannot state the result as \textit{``the asymptotic Fermat's Last Theorem holds over $K$''} since the result applies only to primes $p$ with $p\equiv 2 \pmod{3}$. This restriction comes from Theorem \ref{pomey_case}.
   \end{remark}
   
   Observe that in this result, we do not impose conditions on the S-units solutions over $K$, as in \cite[Theorem 3]{freitas2015asymptotic}. The only conditions we require are about the ramification of certain primes and the class number of $K(\sqrt{-3})$. A similar result, depending only on the class number of $K$ and the ramification of certain primes, is proved, using other techniques, for the equations $x^2+dy^6=z^p$ and $x^4+dy^2=z^p$ in \cite{gpvillagra}.
   
   This presents an advantage since, in general, it is difficult to compute the $S$-unit solutions to the unit equation on $K$. On the other hand, assuming the generalized Riemann hypothesis, it is fast to compute the class number of $K(\sqrt{-3})$. However, it does not present a full alternative to the $S$-unit approach since we only cover half of the primes, i.e. the odd primes such that $p \equiv 2 \pmod3$.

   We can then prove this partial asymptotic result for some particular number fields. One field of interest is     $\QQ(\zeta_{3^n})^+$ ,  the maximal totally real subfield of the cyclotomic extension $\QQ(\zeta_{3^n})$, for which no general result is proven. We are able to prove the following result. 

   \begin{thm}
   \label{cubic_cyclo}
   Let $n$ be a positive integer. Then there exists a bound $C$, depending only on $n$, such that if $p$ is an odd prime with $p \equiv 2 \pmod{3}$ and $p>C$, then there are no non-trivial primitive solutions on $\QQ(\zeta_{3^n})^+$ to the equation $x^p+y^p+z^p=0$.
   \end{thm}

  The article is organized as follows. In section \ref{3|abc} we analyze the case $3|abc$ , where $(a,b,c)$ is a putative solution to the equation \ref{fermat:eq}, and prove Theorem \ref{3midabc}. In order to do so, we use modularity results over number fields, so that we can associate a space of Hilbert modular forms to our Frey curve and then, using that $3|abc$, finally arrive at a contradiction. This is the part where the bound $C_K$ appears. This bound will only depend on the field $K$. In section \ref{3nmidabc} we analyze the case $3\nmid abc$, and prove Theorem \ref{divpomey} and Theorem $\ref{pomey_case}$. Finally, in section \ref{ejemplos} we present some examples of fields that satisfy Assumption~\ref{assumption} and prove Theorem \ref{cubic_cyclo}.

    	
    	
    	
    	
     
    \section{The case $3|xyz$} \label{3|abc}
    
    Let $K$ be a totally real number field, and let $(x,y,z)= (a,b,c) \in \mathcal{O}_K^3$ be a putative solution to the equation (\ref{fermat:eq}). Suppose $3|abc$. As in the proof of the classical Fermat's Last Theorem, we associate to this solution  the Hellegouarch-Frey elliptic curve
    \begin{equation}
    \label{frey-curve}
    E_{(a,b,c)}: y^2 = x(x-a^p)(x+b^p)
    \end{equation}
    defined over $K$. Let $\overline{\rho}_{E,p}$ denote its mod $p$ Galois representation. 

    Let us define 
    \begin{align*}
    S_K = \{\mathfrak{P} : \mathfrak{P} \ \text{is a prime of } K \ \text{above} \ 2\}.
    \end{align*} 
   
    \begin{lemma}
    For all primes $\qq \not\in S_K$, the model of $E_{(a,b,c)}$ given in \eqref{frey-curve} is minimal, semistable and satisfies $p | v_{\qq}(\Delta_{E_{(a,b,c)}})$. The representation $\Bar{\rho}_{E_{(a,b,c)},p}$ is odd, its determinant is the cyclotomic character and is finite flat at every prime $\pp$ of $K$ dividing $p$. Moreover, its conductor is
    \begin{equation}
    \label{conductor}
    \mathcal{N}_{E_{(a,b,c)}} = \prod_{\mathfrak{P} \in S_K} \mathfrak{P}^{r_{\beta}} 
    \end{equation}
    where $0 \leq r_{\mathfrak{P}} \leq 2+6 v_{\mathfrak{P}}(2)$.
    \end{lemma}

    \begin{proof}
    The discriminant of $E_{(a,b,c)}$ is equal to $\Delta_{E_{(a,b,c)}} = 16 \cdot (abc)^{2p}$, then $p|v_{\qq}(\Delta_{E_{(a,b,c)}})$. From \cite[Lemma3.3]{freitas2015asymptotic} we obtain that the model of $E_{(a,b,c)}$ is minimal, semistable and that its conductor is the one expressed in \eqref{conductor}, where $0 \leq r_{\mathfrak{P}} \leq 2+6 v_{\mathfrak{P}}$. The determinant and oddness is a well-known consequence of the theory of the Weil pairing on $E_{(a,b,c)}[p]$. Finally, it follows from \cite{Serre87} that $\Bar{\rho}_{E_{(a,b,c)},p}$ is unramified at $\qq$ if $\qq \nmid p$, and is finite flat at every $\pp$ dividing $p$. 
    \end{proof}

    \subsection{Modularity}
    
    Let $E$ be an elliptic curve defined over a totally real field $K$. We say that $E$ is modular if there exists an isomorphism of Galois representations
    \begin{align}
    \rho_{E,p} \sim \rho_{\ff, \Bar{\omega}},
    \end{align}
    where $\ff$ is a Hilbert cuspidal Hecke eigenform over $K$ of parallel weight $2$ with coefficient field $K_f$ and $\Bar{\omega}$ is a prime in $K_f$ above $p$.

    In (\cite[Theorem 5]{freitas2015elliptic}), Freitas, Le Hung, and Siksek proved the following result about the modularity of elliptic curves.
    
    \begin{thm}
    	Let $K$ be a totally real field. All but finitely many $\overline{K}$-isomorphism classes of elliptic curves over $K$ are modular.
    \end{thm}
    
    In particular, it implies the following corollary.
    
    \begin{coro}
    \label{modular}
    	Let $K$ be a totally real field. There exist some constant $A_K$ depending only on $K$, such that for any non-trivial solution $(a,b,c)$ of the Fermat equation (\ref{fermat:eq}) with prime exponent $p>A_K$, the Frey curve $E_{(a,b,c)}$ given by \eqref{frey-curve} is modular.
    \end{coro}
    
    \begin{proof}
    	See \cite[Corollary 2.1]{freitas2015asymptotic}.
    \end{proof}


    Therefore, if $p>A_K$, there exists a Hilbert cuspidal eigenform $\frak{f}$ over $K$ of parallel weight $2$ and level $\mathcal{N}_{E_{(a,b,c)}}$ such that
    \begin{align}
    \rho_{E,p} \sim \rho_{\ff, \Bar{\omega}},
    \end{align}

    \subsection{Irreducibility}

    To obtain irreducibility of the residual Galois representation $\bar{\rho}_{E_{(a,b,c)},p}$ we use the following result 

    \begin{thm} (\cite[Theorem 6]{anni2016modular})
    \label{irreducible}
    Let $K$ be a totally real field. There is an effectively computable constant $B_K$ such that for a prime $\ell > B_K$, and for an elliptic curve $E/K$ semistable at all $\lambda | \ell$, the mod $\ell$ representation $\Bar{\rho}_{E,\ell}:  G_K \rightarrow \GL_2(\FF_{\ell})$ is irreducible.
    \end{thm}

\begin{remark}
As it is pointed out in \cite{anni2016modular} this result is a restatement of \cite[Theorem 2]{freitas2015criteria}, replacing $K$ with its Galois closure.
\end{remark}
    
Then, since $E_{(a,b,c)}$ is semistable away from $2$ we are under the hypothesis of a lowering the level result. 

\subsection{Level lowering} 

We use the following result from Freitas and Siksek.

\begin{thm} (\cite[Theorem 7]{freitas2015asymptotic})
\label{lowering}
Let $K$ be a totally real field and $E/K$ an elliptic curve of conductor $\mathfrak{N}_E$. Let $p$ be a rational prime. For a prime ideal $\qq$ of $K$, denote by $\Delta_{\qq}$ the discriminant of a local minimal model of $E$ at $\qq$. Let

    \begin{align}
    \mathfrak{M}_p = \prod_{\substack{\qq || \mathfrak{N}_E \\
    p|v_{\qq}(\Delta_{\qq})}} \qq, \ \ \ \mathfrak{N}_p = \frac{\mathfrak{N}_E}{\mathfrak{M}_p}.
    \end{align}
    
Suppose the following conditions hold: 
\begin{itemize}
\item $p\geq 5$, the ramification index $e(\pp|p)<p-1$, and $\QQ(\zeta_p)^+ \notin K$,
\item $E$ is modular,
\item $\Bar{\rho}_{E,p}$ is irreducible,
\item $E$ is semistable at all $\qq|p$,
\item $p|v_{\pp}(\Delta_{\qq})$ for all $\pp|p$.
\end{itemize}
Then, there is a Hilbert eigenform $\ff$ of parallel weight $2$ that is new at level $\mathfrak{N}_p$ and some prime $\Bar{\omega}$ of $K_f$ above $p$ such that $\Bar{\rho}_{E,p} \sim \Bar{\rho}_{f,\Bar{\omega}}$.
\end{thm}

Let us fix $K$ a totally real field. Let $D_K$ a bound such that if $p>D_K$, then we can choose $\pp|p$ with $e(\pp|p)<p-1$, and $\QQ(\zeta_p)^+ \notin K$. If $p>A_K$ then $E_{(a,b,c)}$ is modular by Corollary (\ref{modular}). If $p>B_K$ then $\Bar{\rho}_{E_{(a,b,c)},p}$ is irreducible by Theorem (\ref{irreducible}). Finally, since $E_{(a,b,c)}$ is semistable away from $2$, we conclude that $E_{(a,b,c)}$ satisfies all the hypothesis of Theorem (\ref{lowering}). Therefore there exists a Hilbert eigenform $\ff$ of parallel weight $2$ that is new at level $\prod_{\mathfrak{P} \in S_K} \mathfrak{P}^{r_{\beta}}$ such that
    \begin{align*}
    \bar{\rho}_{E,p} \sim \bar{\rho}_{\ff, \Bar{\omega}},
    \end{align*}

   Fix  a prime ideal $\pp_3$ in $K$ dividing $3$. Since we are in the case that $3|abc$, then $\pp_3|\mathfrak{N}_{E_{(a,b,c)}}$. Let $f= f(\pp_3|3)$. Therefore, since we are in a level lowering case at $\pp_3$ we have that
   \begin{align}
   a_{\pp_3}(\ff) \equiv \pm (3^f + 1) \pmod{p}
   \end{align}
   since this is the necessary condition for Steinberg level raising.

   On the other hand, because of the Hasse-Weil bound, we have that

   \begin{align}
   \label{ineq}
   |a_{\pp_3}(\ff)| \leq 2 \cdot 3^{f/2} .
   \end{align}
   And the same bound applies to all the Galois conjugates of $a_{\pp_3}(\ff)$. In particular, $a_{\pp_3}(\ff) \neq \pm (3^f + 1)$. Indeed, if $a_{\pp_3}(\ff) = \pm (3^f + 1)$ then $a_{\pp_3}(\ff)^2 = (3^f +1)^2 = (3^{f})^2 + 2 \cdot 3^f +1 > 4 \cdot 3^f +1$ which contradicts \eqref{ineq}. From this, we obtain a bound $F_K$ for $p$, since it has to divide the norm of the non-zero algebraic integer $a_{\pp_3}(\ff) \pm (3^f + 1)$. Observe that the bound depends only on $K$ because the form $\ff$ belongs to the space of Hilbert cuspforms over $K$ of parallel weight $2$ and level dividing a power of $2$, depending only on $K$. In particular, the degree of the algebraic integer $a_{\pp_3}(\ff)$ is bounded by the dimension of this space of forms, which depends only on $K$.\\

   Then, we arrive at a contradiction if $p> F_K$.

    

       \begin{proof}[Proof of Theorem \ref{3midabc}]

      It follows from the previous analysis and taking      \begin{equation*}
    C_K = \max\{D_K, A_K, B_K, F_K \}.
    \end{equation*}
\end{proof}

    \section{The case $3 \nmid xyz$} \label{3nmidabc}
    
    \subsection{Introduction to Pomey proof} \label{mimic_pomey}

    We want to analyze now what happens when $3\nmid xyz$. Massouti\'e \cite{massoutie1931dernier} and Pomey \cite{pomey1931nouvelles} proved in $1931$ the following result:
    
    \begin{thm}
    	\label{thm:pomey}
    If $p$ is an odd prime such that $p \equiv 2 \pmod 3$ and $x,y,z$ are non-zero integers such that $x^p+y^p+z^p=0$, then $3$ divides $x,y,$ or $z$.
    \end{thm}
    
    Observe that this result, in particular, proves that there is no non-trivial primitive solution to the equation $x^p + y^p +z^p =0$ regarding $3 \nmid xyz$ and $p$ odd such that $p \equiv 2 \pmod3$, which is the result we are looking for over $K$.
    \begin{itemize}
    	\item \textbf{Question: Does a similar result hold in other fields?} 
    \end{itemize}
    
    
    We mimic Pomey's proof of Theorem \ref{thm:pomey} to see which hypothesis we need to add to $K$ to obtain this result for $K$.
    
    Let $(x_1,x_2,x_3)$ a solution for the equation \eqref{fermat:eq}. Suppose $3 \nmid x_1 x_2 x_3$, then $x_1,x_2,x_3$ are of the form $x_i = \varepsilon_i + 3\lambda_i$ with $\varepsilon_i = \pm 1$. Here, the congruence is in a localized field $\mathcal{O}_K/\pp_3$,  where $\pp_3$ is a prime of $K$ above $3$. Since we want $\mathcal{O}_K/\pp_3$ to be equal to $\FF_3$, we need $T_3 \neq \emptyset$, and this is the reason for the first item of Assumption~\ref{assumption}. Then, we assume $T_3 \neq \emptyset$ and analyze the solutions in $\mathcal{O}_K/\pp_3 = \FF_3$. Since $x_1^p + x_2^p+x_3^p=0$, we have that $\varepsilon_1=\varepsilon_2=\varepsilon_3$. Therefore $3|x_i^2-x_jx_k$. Let define 
    
    \begin{equation}
    	\label{eq:1}
    	P := \frac{x_i^{2p}-x_j^p x_k^p}{x_i^2 - x_j x_k}
    \end{equation}
    
    Observe that
    
    \begin{align*}
    	P = p(x_i^2 x_j x_k)^{\frac{p-1}{2}} + m (x_i^2 - x_j x_k)^2
    \end{align*}
    where
    \begin{align*}
    m=\sum_{r=0}^{p-1} c_r x_i^{2p-4-2r}x_j^r x_k^r, 	\ \ \ c_r = -c_{p-2-r}=r+1 \ \text{for} \ 0\leq r \leq \frac{p-3}{2} 
    \end{align*}

    Since $3|x_i^2-x_j x_k$ and $\varepsilon_j=\varepsilon_k$, 
    \begin{equation}
    	\label{eq:P}
    	P \equiv p \pmod3.
    \end{equation}
     
     On the other hand
    
    \begin{align*}
    	4 (x_i^{2p}-x_j^p x_k^p) = 3(x_j^p+x_k^p)^2 + (x_j^p-x_k^p)^2.
    \end{align*}
    
    Observe that the right hand side is of the form $a^2+3b^2$, with $a,b \in \mathcal{O}_{K(\sqrt{-3})}$. 
    
    In Pomey's proof, he concluded that any positive integer dividing $(x_i^{2p}-x_j^p x_k^p)$, and therefore $P$, will be of the form $a^2+3b^2$. This gives a contradiction since $p\equiv 2 \pmod 3$.

   \subsection{Generalization of Pomey result} \label{general_of_pomey}
   
   Let $K$ be a totally real field and let $x,y \in \mathcal{O}_K$ be coprime integers. Suppose now that $d$ is totally positive and $d| N \coloneqq x^2 + 3y^2$. Under which circumstances can we conclude that $d$ is of the form $a^2+3b^2$?
   
   Observe that $a^2+3b^2 = N(a+b\sqrt{-3})$. Then, $d$ is of the form $a^2 + 3b^2$ if and only if $d = N(\beta)$ for some $\beta \in \mathcal{O}_{K(\sqrt{-3})}$.
   
   Let us denote $\alpha = x+y\sqrt{-3}$ and let
   \begin{align*}
   \alpha \cdot \mathcal{O}_{K(\sqrt{-3})} = \prod_{i=1}^{r} \pp_i^{e_i}
   \end{align*}
   be the prime ideal factorization of $\alpha$ in $\mathcal{O}_{K(\sqrt{-3})}$.

   Since $\pp_i$ are split primes, then $N_{K(\sqrt{-3})/K}(\pp_i)$ are prime ideals. Also, since $x$ and $y$ are coprime integers, the prime ideals $N(\pp_i)$ are coprime to each other. Therefore
   
   \begin{align*}
   	N \cdot \mathcal{O}_K = \prod_{i=1}^{r} N(\pp_i)^{e_i}
   \end{align*}
   is the prime ideal decomposition of $N$ in $K$. Since $d|N$, let
      \begin{align*}
   	d \cdot \mathcal{O}_K = \prod_{i=1}^{r} N(\pp_i)^{e_i'}
   \end{align*}
   be the prime decomposition of $d$ in $K$ with $e_i'$ integers such that $0\leq e_i' \leq e_i$.
   
   Let $t \coloneqq h(K(\sqrt{-3}))$. Therefore, $\pp_i^t = (\alpha_i)$ for some $\alpha_i \in \mathcal{O}_{K(\sqrt{-3})}$. Thus,
   \begin{align*}
   	(d \cdot \mathcal{O}_K)^t &= \prod_{i=1}^{r}  N(\pp_i^t)^{e_i'}\\
   	                          &= \prod_{i=1}^{r} N((\alpha_i^{e_i'}))\\
   	                          &= N((\prod_{i=1}^{r} \alpha_i^{e_i})).
   \end{align*}
   
   Therefore, there exist a unit $u \in \mathcal{O}_K^*$ such that 
   
   \begin{align*}
   	d^t \cdot u = N(\beta)
   \end{align*}
   for some $\beta \in \mathcal{O}_{K(\sqrt{-3})}$
   
   Therefore, $d^t$ is of the form $a^2+3b^2$ if and only if $u = N(\gamma)$ for some $\gamma \in \mathcal{O}_{K(\sqrt{-3})}$.
   
   In particular, if $u=v^2$ for some $v$, then $d^t$ is of that form.
   
   Let us return to the Pomey proof we discussed in the previous subsection. We had $P$ defined in \eqref{eq:1}, which is totally positive. Also, $P$ divides a number of the form $x^2+3y^2$. Then, by the analysis we made, 
   
   \begin{equation}
   	\label{eq:unit}
   	   P^t = u \cdot N(\beta) 
   \end{equation}
   for some $\beta \in \mathcal{O}_{K(\sqrt{-3})}$. Suppose for a moment that $u = N(\gamma)$. Therefore, $P^t$ is of the form $a^2+3b^2$. In particular
   \begin{align*}
   	P^t \equiv \square\pmod 3.
   \end{align*}
   Therefore, using \eqref{eq:P}, we have that 
      \begin{align*}
   	p^t \equiv \square\pmod 3
   \end{align*}
   which is a contradiction if $t$ is odd because $p \equiv 2 \pmod 3$.

  In this last part, we use the condition (3) of Assumption~\ref{assumption}.
   
   \begin{remark}
   	Since $P$ and $N(\beta)$ are totally positive, the unit $u$ will always be totally positive.
   \end{remark}
   
   What we want to see now is, under which hypotheses, the unit $u$ defined in \eqref{eq:unit}, will be a norm of some element in $\mathcal{O}_{K(\sqrt{-3})}$.   
   Let us analyze the following theorem, which will give us information about when this unit $u$ will be a square, and therefore the norm of some integer.
   
   \begin{lemma} 
   	\label{thm:1}
   	Let $K$ be a totally real number field with an odd narrow class number. Then a unit $u$ is a square if and only if it is totally positive.
   \end{lemma}
   
   \begin{proof}
   	Let $C_K$ be the class group of $K$, $C_K^+$ the narrow class group of $K$, $U_K$ the group of units, $U_K^+$ the totally real units, and $K^+$ the totally positive elements of $K$. Then we have the following exact sequence
    \begin{align*}
    1 \rightarrow U_K/U_K^+ \rightarrow K^{\times}/K^{+} \rightarrow C_K^{+} \rightarrow C_K \rightarrow 1
    \end{align*}
    Then,
    \begin{align*}
    |C_K^+| = h \frac{[K:K^+]}{[U_K:U_K^+]} = h \frac{2^r}{[U_K:U_K^+]}
    \end{align*}
    where $h$ is the class group number of $K$ and $r$ its number of real immersions. Since $|C_K^+|$ was odd, $[U_K:U_K^+]=2^r$ and therefore $|C_K^+|=h$. Then, $C_K = C_K^+$. Clearly, if $u$ is a square, then $u$ is totally positive. Suppose now that $u$ is totally positive, that is $u \in U^+$. By the Dirichlet Unit Theorem 
    \begin{align*}
    U = \{\pm 1\} \times \ZZ^{r+s-1}
    \end{align*}
    where $r$ and $s$ are the numbers of real and complex embeddings, respectively. Since $K$ is totally real, $s=0$ and $r=[K:\QQ]$. Then,
    \begin{align*}
    U_K/U_K^2 \cong \{\pm 1\} \times (\ZZ^{r-1})/(\ZZ^{r-1})^2.
    \end{align*}
    Therefore,
    \begin{align*}
    [U_K:U_K^2] = 2 \cdot 2^{r-1} = 2^r
    \end{align*}
    Then $2^r=[U_K:U_K^2]=[U_K:U_K^+][U_K^+:U_K^2]= 2^r[U_K^+:U_K^2]$, and we can conclude that $U_K^+ = U_K^2$
   \end{proof}

   In particular, this implies that if the narrow class number of $K$ is odd, $u$ is a norm, and we are done. However, this hypothesis, jointly with the hypothesis of $h(K(\sqrt{-3}))$ being odd, are really strong assumptions. Then, we will want to deduce a similar result, but with less strong hypotheses, and using hardly that $h(K(\sqrt{-3}))$ is odd. 
   
   First, we will state the following Lemma.
   
   \begin{lemma}
   	\label{class-field-1}
   	Let $K$ be a totally real field. Then $h(K)|h(K(\sqrt{-3}))$. Furthermore, if there exist a prime ideal in $K$ above $3$ ramifying in $K(\sqrt{-3})$ then $h^+(K)|h(K(\sqrt{-3}))$
   \end{lemma}
   \begin{proof}
   	This follows directly from class field theory and the fact that $K(\sqrt{-3})/K$ is a ramified extension since it is a CM field extension.
   \end{proof}
   
   \begin{remark}
   \label{ramification_3}
   	Observe that 
   	  \begin{align*}
   		\mathcal{D}_{K(\sqrt{-3})/K} =\frac{ - (4) \cdot (3)}{(\frak{s} \frak{t})^2}
   	\end{align*} 
   	where $(4)$ and $(3)$ denote their prime factorization in $K$, $\frak{s}$ is the greatest ideal in $K$ such that $\frak{s}^2|(3)$ and $\frak{t}$ the greatest ideal in $K$ such that $\frak{t}^2|2$ and $-3$ is a square $\pmod{\frak{s}^2\frak{t}^2}$. 
   	Then, to guarantee having a prime ideal in $K$ above $3$ ramifying in $K(\sqrt{-3})$, it is enough to have a prime ideal above $3$ of odd ramification index. Here is where we need the condition (2) of Assumption~\ref{assumption}.
   \end{remark}

   Then we have proved the following, 
  
   \begin{proof}[Proof of Theorem \ref{divpomey}]

    Since $V_3 \neq \emptyset$, there exist a prime ideal in $K$ above $3$ ramifying in $K(\sqrt{-3})$. Then $h^+(K)$ is odd by Lemma \ref{class-field-1}. Therefore, the unit $u$ described at the beginning of the subsection \ref{general_of_pomey}, i.e. the unit such that $d^t \cdot u = N(\beta)$, is a square by Lemma \ref{thm:1}. Therefore, $d^t$ is of the form $x^2+3y^2$. 
\end{proof}
   
      \begin{proof}[Proof of Theorem \ref{pomey_case}]

    Follows directly from Theorem \ref{divpomey} and the analysis made at the subsection \ref{mimic_pomey}.
\end{proof}

  Finally, from Theorem \ref{pomey_case} and Theorem \ref{3midabc} we conclude Theorem \ref{thm:main}.
   
   

\section{Examples of fields K} \label{ejemplos}
   
We want to give an example of totally real fields $K$ satisfying Assumption~\ref{assumption}. A natural question is whether there exists a quadratic extension satisfying the assumption. 
    
    \begin{prop}
    	\label{quadratic}
     Let $p$ be an odd prime. Then there are no quadratic totally real extensions $K$ of $\QQ$ such that $h(K(\sqrt{-p}))$ is odd and such that $p$ splits in $K$.
    \end{prop}
    
    \begin{proof}
    Suppose there exist such a quadratic extension $K:=\QQ(\sqrt{d})$ and that $\pp$ splits as $\pp_1 \pp_2$ in $K$. Then, $K(\sqrt{-p})$ ramifies at $\pp_1$ and $\pp_2$ and probably at the primes above $2$ in $K$. By Lemma \ref{class-field-1}, we also have that $h(K)$ is odd. Let $\mathcal{A}_{K(\sqrt{-p})/K}$ be the number of ambiguous classes of $K(\sqrt{-p})/K$ which are in the $2$-class group $K(\sqrt{-p})$. Then using the \textit{ambiguous class number formula} of Chevalley (see \cite[Chapter 13, Lemma 4.1]{lang2012cyclotomic} and \cite[Section 3]{mccall1995imaginary}) we obtain that
    \begin{align*}
    	\mathcal{A}_{K(\sqrt{-p})/K} = \frac{2^{1+\gamma} }{[\mathcal{O}_K^{\times}:\mathcal{O}_K^{\times} \cap \text{Norm}(K(\sqrt{-3})^{\times})]}
    \end{align*}
    where $\gamma$ is the number of prime ideals of $K$ ramifying in $K(\sqrt{-p})$. Since $K(\sqrt{-p})$ ramifies at $\pp_1$ and $\pp_2$, we obtain that $\gamma \geq 2$. Since $K$ is totally real and $-1$ is not a norm we have that $[\mathcal{O}_K^{\times}:\mathcal{O}_K^{\times} \cap  \text{Norm}(K(\sqrt{-p})^{\times})]= 2^e$ with with $1 \leq e \leq 2$. Then, $2|\mathcal{A}_{K(\sqrt{-p})/K}$. However, this is a contradiction since $\text{Cl}^{(2)}(K(\sqrt{-p})) = \{e\}$ because $h(K(\sqrt{-p}))$ was odd. Then we conclude that there not exist a quadratic extension of $\QQ$ satisfying what we wanted.
    \end{proof}
    
    Therefore, it implies in particular the following Corollary.
    \begin{coro}
    There are no quadratic totally real extensions of $\QQ$ satisfying  Assumption~\ref{assumption}.
    \end{coro}
    \begin{proof}
    This follows from Proposition \ref{quadratic} and the fact that satisfying Assumption~\ref{assumption} in a quadratic field is equivalent to asking that $3$ splits.
    \end{proof}

For a general totally real field we have the following result. 
        
    	\begin{prop}
    		Let $p$ be an odd prime. Let $K$ be a totally real field with $h(K(\sqrt{-p}))$ odd and such that $V_p \neq \emptyset$. 
            Then $V_p = \{\pp\}$ with $\pp$ a prime ideal of $K$ above $p$. In particular,  the only odd prime ideal ramifying in $K(\sqrt{-p})/K$ is $\pp$. 
    	\end{prop}
    	
    	\begin{proof}
                Suppose there exists a prime ideal $\pp' \in V_p$ different from $\pp$. In particular, it ramifies in $K(\sqrt{-p})/K$. Then, using the Chevalley formula for ambiguous class numbers, we obtain that
                  \begin{equation}
                  \label{eq:ambiguous}
    	    \mathcal{A}_{K(\sqrt{-p})/K} = 
                 \frac{2^{\gamma + r + s -1} }{[\mathcal{O}_K^{\times}:\mathcal{O}_K^{\times} \cap \text{Norm}(K(\sqrt{-3})^{\times})]}
                 \end{equation}
                 where $\gamma$ is the number of finite prime ideals of $K$ ramifying in $K(\sqrt{-p})$, and $r$ and $s$ are the real and complex embeddings of $K$ respectively. Since $K$ is totally real, the primitive roots of unity in there are $\{\pm 1\}$. Then, using Dirichlet's unit theorem we obtain that $[\mathcal{O}_K^{\times}:\mathcal{O}_K^{\times} \cap \text{Norm}(K(\sqrt{-3})^{\times})] = 2^k$, with $0\leq k \leq r+s$. Therefore, from \eqref{eq:ambiguous} we have that $\mathcal{A}_{K(\sqrt{-p})/K} = 2^e$ with $\gamma - 1 \leq e$. Since we have, at least two prime ideals $\pp$ and $\pp'$ of $K$ ramifying in $K(\sqrt{-p})$ we deduce that $\gamma \geq 2$. Therefore $\gamma -1 \geq 1$ and thus $2|\mathcal{A}_{K(\sqrt{-p})/K}$, arriving to a contradiction since $h(K(\sqrt{-p}))$ was odd.
    	\end{proof}

\begin{remark}
\label{options}
If a totally real field satisfies the hypothesis of Theorem \ref{thm:main}, it has to satisfy then that $V_3=\{\pp_3\}$. The  possibilities for the ramification of $3$ are the following: 
        \begin{enumerate}
        \item $3$ totally ramifies in $K$ and $[K:\QQ]$ is odd,
        \item $3$ splits as $\pp_0^k \cdot \prod_{i=1}^n \pp_i^{2k_i}$ where $k$ is odd and $f(\pp_{i_0}|p)=1$ for some $\pp_{i_0}$.
        \end{enumerate}
        
        In particular, in the second case $K$ cannot be a Galois extension.
\end{remark}

     \subsection{Cubic extensions}   

     In \cite{alvarado2021robust}, while developing an algorithm to compute the $S$-unit solutions, they listed in Table 8.1 the totally real cubic extensions $K$ in which $2$ totally ramifies and $|\Delta_K| \leq 2000$. Then they verify that the $S$-unit solutions satisfy either condition (A) or (B) of \cite[Theorem 3]{freitas2015asymptotic}.

     In our case, we do not have to deal with $S$-unit equation, so the computation is straightforward since we only have to compute $h(K(\sqrt{-3}))$ and $3 \cdot \mathcal{O}_K$. Let $\mathscr{K}_X$ denote the set of totally real number fields satisfying Assumption~\ref{assumption} with $h(K(\sqrt{-3}))$ odd and which have absolute discriminant $\Delta_K$ satisfying $|\Delta_K|\leq X$. Table \ref{table:1} lists all the fields of $\mathscr{K}_X$ for $X=2000$. In particular, it follows from Theorem \ref{thm:main} that the Asymptotic Fermat Last Theorem, for exponent $p \equiv 1 \pmod{6}$, holds for these cubic fields listed.

     \begin{table}[h!]
    \centering
 \begin{tabular}{llll}
        \toprule
        $f_K$ & $|\Delta_K|$ & $h(K(\sqrt{-3})$ & Ramification of $3$\\
        \midrule
        $x^3-3x-1$ & $81$ & $1$ & $\pp^3$\\
        $x^3-x^2-4x+1$ & $321$ & $1$ & $\pp_1 \pp_2^2$\\
        $x^3-x^2-5x+3$ & $564$ & $1$ & $\pp_1 \pp_2^2$\\
        $x^3-6x-3$ & $621$ & $1$ & $\pp^3$\\
        $x^3-6x-2$ & $756$ & $1$ & $\pp^3$\\
        $x^3-6x-1$ & $837$ & $1$ & $\pp^3$\\
        $x^3-x^2-6x+2$ & $993$ & $1$ & $\pp_1 \pp_2^2$\\
        $x^3-x^2-9x+12$ & $1101$ & $1$ & $\pp_1 \pp_2^2$\\
        $x^3-x^2-8x-3$ & $1425$ & $1$ & $\pp_1 \pp_2^2$\\
        $x^3-x^2-7x+1$ & $1524$ & $1$ & $\pp_1 \pp_2^2$\\
        $x^3-12x-14$ & $1620$ & $1$ & $\pp^3$\\
        $x^3-9x-6$ & $1944$ & $1$ & $\pp^3$\\
        \bottomrule
        \end{tabular}
\caption{Fields in $\mathscr{K}_{2000}$ and how $3$ splits.}
\label{table:1}
\end{table}

\begin{remark}
Observe that $3$ splits in the way explained in Remark \ref{options}. Case $1$ of the Table \ref{table:1} corresponds to the totally real subfield of $\QQ(\zeta_9)$. In the subsection \ref{cyclotomic}, we deal deeply with these cases.
\end{remark}

 \subsection{Cyclotomic extensions} \label{cyclotomic}

Let $L$ be a cyclotomic extension. Suppose $L=\QQ(\zeta_{p^k})$ for $p$ an odd prime number and $k$ a positive integer. We first would like to find examples that satisfy the hypothesis about the splitting of $3$ of Theorem \ref{thm:main}. Let $K$ be the totally real subfield of $L$. Since $K$ is a Galois extension, $3$ has to ramify in $K$ as it is explained in \ref{options}. Then, the only examples that may satisfy the desired hypothesis are the fields $\QQ(\zeta_{3^n})^+$.The following result guarantees us the odd parity of the class number of these fields.







\begin{prop}
\label{prop-parity}
(\cite[Proposition 3]{HORIE_2002}) Let $\ell$ be a prime congruent to $2$ or $5$ modulo $9$. Then, for any positive integer $n$, the class number of the cyclotomic field $\QQ(\zeta_{3^n})$ is relatively prime to $\ell$.
\end{prop}

Then, using this latter result we are able to prove that the main theorem \ref{thm:main} works on $\QQ(\zeta_{3^n})^+$ as it is stated in \ref{cubic_cyclo}.

\begin{proof}[Proof of Theorem \ref{cubic_cyclo}]
Let us denote $K=\QQ(\zeta_{3^n})^+$. Observe that $K(\sqrt{-3})= \QQ(\zeta_{3^n})$, then $h(K(\sqrt{-3}))$ is odd by Proposition \ref{prop-parity}. Since $[K:\QQ]=3^{n-1}$ we have that $S_3=T_3=V_3=\{\pp_3\}$ with $\pp_3$ the only prime above $3$ in $K$. Then the result follows from Theorem \ref{thm:main}.
\end{proof}


In \cite[Theorem 4]{freitas2020asymptotic}, Freitas, Kraus, and Siksek proved that under the condition of $2$ being inert in the maximal subfield of $\QQ(\zeta_{q^r})^+$ of $q$-power degree, the asymptotic Fermat's Last Theorem holds for such a field as long as $q\geq 5$. As it is stated in \cite[Remark 3.3]{freitas2020asymptotic}, the proof of Theorem 4 fails for $q=3$ since there are unit solutions to the equation 
\begin{equation}
\label{eq:S-unit}
\lambda + \mu = 1, \ \ \lambda, \mu \in \mathcal{O}_K^+.
\end{equation}

However, for our result, we do not have to deal with the $S$-unit equation, which allows us to cover the case $q=3$, at least for half of the prime exponents. Observe that another important difference with the case $q > 3$ is that we obtain a result for $q = 3$ over the full extension $\QQ(\zeta_{3^r})^+$. In \cite[Theorem 1.2]{chen2024fermat}, Chen et al. proved a similar result for the field $\QQ(\sqrt{5})$. They established Fermat's Last Theorem over $\QQ(\sqrt{5}) = \QQ(\zeta_5)^+$ for all prime exponents $p$ satisfying $p \equiv 5, 7 \pmod{8}$ or $p \equiv 19, 41 \pmod{48}$. Note that this is not an asymptotic result, since it holds for all such primes $p$. Aside from this result, for $q > 3$, no results on Fermat's Last Theorem are known over the full extensions $\QQ(\zeta_{q^r})^+$; the known results hold only over proper subfields of these fields.

Finally, let us observe the following:

\begin{prop}
$2$ is inert in $\QQ(\zeta_{3^n})^+$ for all $n$.
\end{prop}
\begin{proof}
Follows directly from \cite[Lemma 2.1]{freitas2020asymptotic}
\end{proof}

Then, the fields $\QQ(\zeta_{3^r})^+$ are examples of fields in which the asymptotic Fermat's  Last Theorem holds for primes $p$ such that $p \equiv 2 \pmod{3}$, but do not satisfy neither condition (A) nor (B) of \cite[Theorem 3]{freitas2015asymptotic}. Indeed, let $\mathfrak{P}$ be the prime ideal above $2$. Then, (A) does not hold since in this case $T=\emptyset$ and (B) does not hold since there are unit solutions $(\lambda,\mu)$ to the unit Equation (\ref{eq:S-unit}) (see \cite[Lemma 16]{siksek2024curves}). Then $v_{\mathfrak{P}}(\lambda \mu) = 0 \not\equiv 1 \pmod{3}$ where $S=\{\mathfrak{P}\}$.

\bibliographystyle{plain}
\bibliography{biblioppp}
\end{document}